OpenStudy (anonymous):

A town of 10,000 people has been growing by 300 people each year A second town of 30,000 has been shrinking by 200 people each year.  When will the two towns have the same population? i think i may have the correct the answer but i just want to double check with you guys

OpenStudy (anonymous):

10000+300n=30000-200n n=number of years 500n=40000 n=80

OpenStudy (anonymous):

actually is 500n=20000 and not 40000 so n=40.